OpenStudy (anonymous):

Write the equation in logarithmic form. 2^5 = 32 A. log32 = 5 • 2 B. log232 = 5 C. log32 = 5 D. log532 = 2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Hint: \[\Large b^x = y\] converts to \[\Large \log_{b}(y) = x\]

OpenStudy (anonymous):

that confused me a bit

jimthompson5910 (jim_thompson5910):

in your case, b is 2, x is 5 and y is 32

jimthompson5910 (jim_thompson5910):

do the proper replacements

OpenStudy (anonymous):

log32 = 5 times 2??

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

o haha im so confused

jimthompson5910 (jim_thompson5910):

ex: \[\Large 3^4 = 81\] converts to \[\Large \log_{3}(81) = 4\]

OpenStudy (anonymous):

how do i put that 3 in my calculator

jimthompson5910 (jim_thompson5910):

\(\Large 3^4\) = 3 ^ 4

jimthompson5910 (jim_thompson5910):

and there's no need to use your calculator for this problem

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

so what am i doing lol

jimthompson5910 (jim_thompson5910):

you're just going from something like \[\Large 3^4 = 81\] to \[\Large \log_{3}(81) = 4\]

jimthompson5910 (jim_thompson5910):

except I used different numbers

OpenStudy (anonymous):

im 100% LOST

jimthompson5910 (jim_thompson5910):

\[\Large 3^4 = 81\] is in exponential form \[\Large \log_{3}(81) = 4\] is equivalent, but its in logarithmic form

jimthompson5910 (jim_thompson5910):

that's all you're doing: converting from exponential form to logarithmic form

OpenStudy (anonymous):

my calculator wont let me do it error and idk how to do it without it

jimthompson5910 (jim_thompson5910):

you don't need a calculator, all you're doing is rearranging the numbers

jimthompson5910 (jim_thompson5910):

the base becomes the base of the log the exponent goes on the right side the number that was on the left side goes inside the log

jimthompson5910 (jim_thompson5910):

 becomes